Study of Top Quarks

in Hadronic Interactions

Pekka K. Sinervo University of Toronto

Lecture # 1 Discovery of the Top Quark

1. Hadron Colliders and Detectors

2. The Counting Experiments

Lecture # 2 Top Quark Properties

1. Cross section measurements

2. Production and Decay Kinematics

3. Mass and BR Measurement

Lecture # 3 High Luminosity Top Studies

1. Search for New Physics

2. High Statistics studies of Top

3. Future Top Quark studies

1995 SLAC Summer Institute

SM Predictions for Top Quarks

The asymmetry in e^{+} e^{−} → bb̄ scattering implies that b is part of SU(2) doublet

• top quark has standard SU(2) couplings

• CKM matrix predicts dominate decay t → W + b

W^{±} and Z^{∘} masses depend on heaviests fermions

• LEP, SLD measurements on the Z resonance require

Mtop = 179 ± 9{+ 17 \atop −19} GeV/c2

D. Schaile, 1995 pp̄ Topical Conference

Top Quark Phenomenology

Top Quarks decay via the weak force

• Predict ∼ 100% t → W + b

• Lifetime depends on mass

• For heavy top (Mtop > ∼ 100 − 120 GeV/c2), decays before it hadronises

Sensitive to some classes of “new physics”

• Higgs couples strongly to top

• Could decay into a SUSY H+

t → H+ b

|_{⟶}τ+ ντ

• Top quark can be sensitive to SM details – Vts CKM matrix elements

– Constraints on other SM parameters (eg. ALR)

– New physics that couples to t t system

What Do We Already Know About Top?

Searches at e^{+} e^{−} Colliders have ruled out “light” top

• data from LEP requires Mtop > 46 GeV/c2 at 95% CL • Independent of how top decays

Width of W ± boson rules out W + → tb̄

• Mtop > 62 GeV/c2 at 95% CL

F.Abe et al. (CDF), PRL 73, 220 (1994).

Searches at pp̄ colliders at CERN and Fermilab

• searched for a “Standard Model” top

• Mtop > 131 GeV/c2 at 95% CL (D0 Collab).

S. Abachi et al. (D0 / ), PRL 72, 2138 (1994).

Production of Top in Hadron Collisions

Top quark production is mediated by

qq̄, gg → tt̄
qq̄ → tb̄X  and  qq̄ → W + → tb̄ both small

Because top is heavy, need large x for initial partons

• Implies that cross section falls with mass

• t t system approximately at rest – < PT >∼ Mtop/2

Higher order corrections to σ_{t t} have been calculated

• Most recently by Laenen, Smith and Neerven, NP B369, (1992) 543.

• Biggest effects due to “soft gluon emission”

Top Decays and Detection Requirements

SM top quark decays via the weak force

t → W + b

|_{⟶}ud̄ or cs̄ (2/3)

|_{⟶}e+ νe or μ+ νμ or τ+ ντ (1/3)

In proton-antiproton collisions,

This results in several distinct decay signatures

• Six jet final states (36/81)

• Lepton (e± or μ±) + Jet events (24/81)

• Dilepton events with 2 energetic e± and μ∓ (4/81)

• Decay modes with τ’s hard! (∼ 12/81)

Detector Requirements

Detection of these decay modes requires

• Good jet reconstruction (energy and direction)

• Good lepton ID capability (acceptance and efficiency) – > 80 − 85% is reasonable

• Missing energy detection (for neutrinos)

Production/decay properties imply that

• Production is “central” (especially for high mass)

• Most likely initial state is where x1 ∼ x2

• Top quarks are produced with modest PT

⇒ daughters are isotropic in lab frame

Detector requirements boil down to

1. Coverage over |η| < ∼ 2 − 2.5 desirable

2. Good efficiency for partons with PT ≥ 15 − 20 GeV/c 3. Good lepton ID with high efficiency

Can see this from t t MC

• Mtop = 160 GeV/c2 in dilepton channel • Plot observed jet and lepton energies

Background Considerations

The backgrounds are dominated by QCD processes

• Multi-jet final states direct background to 6-jet channel • Jets faking leptons potential background to lepton ID • Many SM EWK physics processes

– Drell-Yan production in e+ e− and μ+ μ− modes

– Z∘ production

– Semi-leptonic bottom and charm quark decays

– QCD W + jet processes

– Di-vector boson final states: W + W −, W + Z∘

Any of these could compromise one of the searches

• Lepton ID must have extremely good jet-rejection

• Kinematics of background events must be used

– For example, can use PT and lepton isolation requirements to reduce and characterise b, c quark production

Background Studies Extremely Important

The Fermilab Collider Experiments

CDF and D0 / designed to study p p interactions

• Collisions with √s = 1.8 TeV created by Tevatron Collider

• Sophisticated general-purpose detectors operated by big collaborations

Main detector features:

• Hermetic Calorimetry

– D0 / has state-of-the-art Ur/Li Ar calorimeter ∗ detection of e±, γ and jets

– CDF has more traditional Pb and Fe calorimeter ∗ Scintillator readout for central region |η| < 1.1

∗ Gas proportional chamber readout for plug/forward – Coverage over |η| < ∼ 4.2

– Good hermiticity for E̸T measurement

– Provides electron ID and jet detection

Resolutions  D0 /  CDF
EM  0.15√E ⊕ 0.004E  0.13√E ⊕ 0.02E
π±, K±  0.50√E  0.80√E
E̸T x,y  1.08 + 0.019 ∑ ET

Detectors (continued)

• Muon chambers outside calorimeter – identify charge tracks that pass through calorimeter

– D0 / has toroidal magnets to measure muon PT ∗ ⇒ σp/p ∼ 0.18p

• Central Tracking Systems

– CDF has three-tiered system in 1.4 T field ∗ Central Track Chamber (CTC) for 28 < R < 130 cm ∗ Vertex Chambers (VTX) in 10 < R < 23 cm

∗ Silicon Vertex Detector (SVX) in 3 < R < 12 cm

∗ ⇒ σp/p ∼ 0.001p (in GeV/c)

– D0 / has nonmagnetic central tracker

∗ Also incorporates a Transition Radiation Detector (TRD) for electron ID

By design, detectors are complementary

• D0 / emphasises calorimetery and muon ID

• CDF emphasises tracking and electron ID

• Reality is that both have strengths and weaknesses

Detector Limitations

CDF’s dirty laundry”

• Muon system – not enough absorber

– coverage somewhat fragmented

• Calorimetry – Not totally hermetic

– Region |η| > 3.6 not well understood

D0 / ’s laundry list:

• Trigger system still being shaken down

• Main ring backgrounds

• Lack of magnetic field reduces redundancy – Calibration more difficult

– Not able to detect secondary vertices

Triggering Issues

Use multi-level trigger to tag events in real-time

• Interaction rate of (50 × 10−27 cm 2) ×(1.5 × 1031 cm −2s−1) = 7.5 × 105 Hz has to be reduced to of order 1-10 Hz

• Most challenging aspect of experiment – Depends critically on all subsystems

– Need efficiency > 75% to be competitive

Triggering for “heavy” top is relatively easy

• Tag high PT (> 15 − 20 GeV/c) lepton(s)

• Missing energy requirements (E̸T > 20 − 25 GeV)

• Multijet topologies

• Trigger efficiencies ≥ 90 − 95% in most cases

1992-95 Collider Run

Tevatron Collider has been running since 1992

• Run 1A (1992-93) provided ∼ 20 pb−1

– CDF scanned ∼ 1 × 1012 collisions in 19.3 pb−1 ∗ Collected 16 × 106 events on tape

– D0 / collected 13.5 pb−1

• Run 1B (1994-95) has produced another ∼ 85 pb−1 – Run expected to continue till end of 1995

– Show results from 50 − 70 pb−1

By far the largest and most sensitive dataset to top

• For a top quark with a mass of 160 GeV/c2

– σt t = 8.2 pb ⇒ 400 produced t t events

– D0 / and CDF would each expect to detect ∗ ∼ 6 dilepton events

∗ ∼ 20 lepton + jet events

Dilepton Top Quark Search

t t decaying into 2 leptons create unique signature

• Two charged leptons from W ± decay

• Imbalance in total energy due to undetected neutrinos • Two jets from bottom quarks

CDF performs search requiring

• Two clean, high PT lepton candidates

• Missing transverse energy E̸T > 25 GeV

• Require two jets with ET > 10 GeV and |η| < 2.0 • Reject specific backgrounds (e.g. Z∘)

Significant dilepton background is Drell-Yan production

• E̸T signature created by mismeasured jet

• Harden E̸T cut to > 50 GeV if correlated with jet

a) Z+1 jet, b) Z+ 2 jet, and c) t t (Mtop = 160 GeV/c2)

Example: CDF Z → τ^{+} τ^{−} Background

The Z → τ^{+} τ^{−} → e^{+} μ^{−}ν_{e}ν̄_{μ}X is largest background for all dilepton channels

• Take sample of observed Z → e+ e− and Z → μ+ μ− events – Gives correct Z kinematics

– Gives correct associated jet production

• Replace each lepton with a τ± – Use MC to decay τ±

– Use detector simulation on τ± decay daughters

Resulting calculation is relatively model-independent

• MC model dependence very small

• Limited by Z statistics

Jet Cut  e^{+} e^{−} and μ^{+} μ^{−}  e^{±}μ^{∓}
Jet E_{T} > 10 GeV  0.17 ± 0.05  0.21 ± 0.05
Jet E_{T} > 15 GeV  0.10 ± 0.03  0.13 ± 0.04

CDF Results for Dilepton Channel

Have measured the rate of events from non-top sources

• Used data to measure most of these rates

• Cross-checked using Z∘ → e+ e− and Z∘ → μ+ μ−

• Consistent with theoretical calculations

Background  eμ (events)  ee, μμ (events)
Z → τ+ τ−  0.21 ± 0.05  0.17 ± 0.05
One or both leptons fake  0.10 ± 0.10  0.13 ± 0.10
bb̄  0.01 ± 0.01  0.02 ± 0.01
W W, W Z and ZZ  0.13 ± 0.06  0.08 ± 0.03
Drell-Yan  0.44 ± 0.28
Total  0.45 ± 0.14  0.84 ± 0.28

• Expect 1.3 ± 0.3 background events

Control of Backgrounds

Evolution of background rates in dileptons instructive

Run  Sensitivity  Bkg σ Comments
Period  (pb−1)  (fb)
1988-89  4  350  Just dilepton requirement
CDF 1A  20  20  More Drell-Yan rejection
CDF Run 1A/B  67  19  Added ≥ 2 jet cut
D0 / 1A  15  72  ≥ 1 jet (for Mtop< 120)
D0 / 1A  15  52  ≥ 2 jets (for Mtop> 120)
D0 / 1A/B  44-56  14  Added HT cut

Have to be careful with background studies

• Data fluctuations can lead to “carving up” of acceptance

• Base cuts on independent samples

• Use MC predictions to guide

What Is Expected From Top

Expected number of t t→ dileptons is determined using MC and full detector simulation

Mtop  σt t  Events
(GeV/c2) (pb) Detected
160  8.2  4.4
170  5.8  3.0
180  4.2  2.4

Breakdown by channel reflects branching ratios

Channel Signal Fraction
eμ  57%
ee  20%
μμ  23%

Mtop of 170 GeV/c2 assumed

Dilepton acceptance as a function of top mass

21

Dielectron Data for Run 1A and 1B

22

DiMuon Data for Run 1A and 1B

23

Electron-Muon Data for Run 1A and 1B

24

Summary of CDF Dilepton Search

Observe a total of 6 candidate events

• Observe 1 μ+ μ− event and 5 eμ events – Additional μ+ μ− event excluded

– Has μ+ μ−γ mass close to Z∘

• Expect 1.3 ± 0.3 events from background

Probability of background fluctuation is 3.3 × 10^{−3}

This means:

• “If experiment is repeated and top doesn’t exist, then fraction of experiments with at least 6 events observed is 1 in 300.”

B Quark Tags in CDF Dilepton Events

Can a posteriori check for evidence of b jets

• Use SVX and SLT algorithms described later – Two events with SVX-SLT “double tags”

– One event with SLT μ tag

• Would expect 0.47 tags if no b quark jets – Would expect 3.6 tags if all 6 are t t

D0 / Results for Dilepton Channel

D0 / applies slightly different P_{T} cuts on leptons

• Goal is to optimise signal to noise – Lower lepton thresholds for eμ channel

– Use HT kinematic cut

HT ≡ ∑_{jets}ET

• Add electron ET to HT for ee and eμ channels

• ≥ 2 jets with ET > 15 GeV and |η| < 2.4

eμ  ee  μμ
Lepton PT  > 15/12 GeV/c  > 20 GeV/c  > 15 GeV/c
E̸T  > 20 GeV  > 25 GeV
HT  > 120 GeV  > 120 GeV  > 100 GeV

• Jet cuts are looser than for CDF

• Lepton thresholds are lower

• Use HT cut to get comparable rejection

Data compared with expectation from

• Background modelled by Z production • Signal is t t ISAJET MC

• Shaded are events that pass all other cuts

Results of D0 / Dilepton Search

Background predicted to be Drell-Yan and Z → τ^{+} τ^{−}.

eμ  e+ e−  μ+ μ−
∫ Ldt (pb−1)  48 ± 6  56 ± 7  44 ± 5
Data (events)  2  0  1
Backgrounds (events)  0.12 ± 0.03  0.28 ± 0.14  0.25 ± 0.04
Mtop = 140 GeV/c2  1.4  1.0  0.5
Mtop = 160 GeV/c2  0.9  0.7  0.4
Mtop = 180 GeV/c2  0.6  0.4  0.2

Observe 3 events and expect 0.65 ± 0.15 background

• Probability of background fluctuation is 0.03

• There is excess but not significant on its own

Summary of Dilepton Results

CDF/D0 / observe nine events

• For CDF, results are very suggestive – Five electron and muon candidates

– Three have b quark jet tags

• Expect a total of 2.0 ± 0.4 events background

Expected rate from top is small and depends on mass

Mtop (GeV/c2) CDF events  D0 / events
160  4.4  2.0
170  3.0  1.5
180  2.4  1.2

Statistically significant observations:

• Probability 1.3 events fluctuate to ≥ 6 events is 3.3 × 10−3

• Probability 0.3 events fluctuate to ≥ 3 events is 3 × 10−2

⇒ Have observed strong evidence for t t

Lepton + Jets Search Philosophy

Decay mode with the greatest sensitivity is “lepton+jets”

• About 30% of t t events decay this way

• Look for one high PT lepton

• Require E̸T > 25 GeV to “tag” neutrino

Problems come from physical backgrounds

• Backgrounds from “fake” leptons, etc. are manageable

• “QCD W + jet” production much larger than t t

• Have to find additional rejection power 1. Place more stringent kinematical requirements 2. Require presence of b quark jet

– “tag” semileptonic b quark decays

– “tag” using large cτ of b decays

∗ This capability is unique to CDF

• Have to maintain good top efficiency

W+Jet Selection

Selection of W candidates straightforward

• Require high PT isolated lepton candidate

• Require significant E̸T

• Require three or more high PT jets

Both experiments have large inclusive W samples

• CDF has 31,000 eνe and 21,000 μνμ candidates

• D0 / has corresponding numbers

Run 1A CDF data shows clear “W” signal:

W + Jet Sample

Require ≥ 3 or 4 jets (from other W and b quarks)

• CDF: ≥ 3 jets with ET > 15 GeV and |η| < 2.0 – finds 203 W+≥ 3 jet events

• D0 / uses somewhat looser jet cuts – ET > 15 GeV (corrected to parton ET), |η| < 2.4 – Places additional kinematic cuts

Most background sources are small:

Background  Number of Events (CDF)
Fake leptons and bb̄  18 ± 9
W W, W Z and ZZ  9 ± 4
Z → ττ, ee, μμ  15 ± 3
Total  42 ± 10

QCD W+jets production is largest expected background

Mtop (GeV/c2) Number of Events (CDF)
120  136 ± 21
140  77 ± 10
160  42 ± 7
180  23 ± 2.4
W+jets  175{+56 \atop −28}

Jet Multiplicity

The D0 / data illustrates the serious S/N problem

• Event distribution versus minimum jet multiplicity • Require jet ET > 15 GeV (corrected)

Event Topology

Can use kinematics to separate t t from QCD background

• t t events are more spherical (aplanarity)

Mij = ∑_{jets} \frac{pipj}{p2}

A ≡ \frac{3}{2} Q1 where Q1 is smallest eigenvalue of M

• t t events have more energetic partons (HT)

HT = ∑_{jets}ET

D0 / has applied both techniques to W+≥ 4 jets

• Require A > 0.05 and HT > 200 GeV

• Estimate W+jet background by scaling jet multiplicities • Observe 8 events and expect 1.9 ± 0.5 from background 36

W+jet Background Scaling

D0 / estimates background in W + ≥ 4 jet sample

• assumes ratio N(l + n jets )/N(l + (n − 1)jets ) constant Berends et al., NP B357, 32 (1991).

• checked with Z+jets, W +jets (n = 2, 3)

• Determine Ntop in W + ≥ 4 jet sample (w/o A, HT) – Apply A, HT cuts, estimate effect on signal – Surviving background is 1.9 ± 0.5 events

Need to know efficiency of A, H_{T} cuts on signal and background

• Checked by varying cuts

• Using Z+jets sample (statistically limited)

Jet Scaling Calculation (cont)

Indications are that scaling works well

• Have relatively small uncertainty on extrapolation

B Tagging Techniques

Use two properties of b quarks to identify them

1. About 20% of b’s decay via

b → e−νeX and b → μ−νμX

• Look for jets with “soft” leptons in them (SLT tags)

• CDF searches for both electrons and muons

• D0 / only uses soft muons

– has higher efficiency for “cascade” decays

b → cX → μνμX′

2. “Long” b lifetime – cτ ∼ 450 microns

• CDF uses Silicon Vertex Detector to identify jets with a secondary vertex

• We call these “SVX” tags

• Expect 42% of t t events to have at least one SVX tag

Idea is to just “count” number of tagged events

Soft Lepton Tagging

CDF identifyies muon or electron from b hadron decay

• Require lepton PT > 2 GeV/c

• Require electron to satisfy standard cuts – Use dEdx measurement in CTC to reject pions

– Require muon candidate to have good quality muon stub

D0 / searches for muons with P_{T} > 4 GeV/c

• Lepton threshold does not effect acceptance significantly • Expects to be able to tag about 20% of t t events

CDF has studied SLT tagging using

• photon conversions for electrons (γ → e+ e−)

• J/ψ → μ+ μ− decays

• generic jets to obtain fake rates – fakes primarily hadron punch-through

– “Fake” tagging rate is ∼ 0.005 − 0.010 tags/track

• Efficiency is 0.152 ± 0.023 for Mtop = 160 GeV/c2

SVX B Tags

SVX tag requirements are

• at least 3 tracks with impact parameter d > 3σ – use 2 tracks with tighter cuts in 2nd pass

• track PT > 2 GeV/c

• the tracks are consistent with coming from a common vertex

Used control samples to measure efficiency, fake rate

• inclusive electrons enriched in b → e+ νeX decays

• B → J/ψX, J/ψK±, J/ψK∗ decays

• “generic” jet sample

SVX B Tagging

Can demonstrate that tags are really finding b’s

• Inclusive e±’s are ∼ 40 − 60% from b → e−νec • Look at jet “away” from e± candidate

• See expected number of tags

SVX B Tagging Backgrounds

Can use rate of tags in QCD jet data to estimate background

• Multiply number of jets by per-jet tagging rate ft – ft ∼ 0.01 and is independent of jet pT

• This includes real pp̄ → bX

• Include MC estimate for W c production

• Does provide a conservative background estimate

More rigorous to separate background sources

• Fake tags estimated from rate of −ve decay length tags – Turns out to be ∼ 0.003 tags per jet

• Tags from real heavy flavour (W bb̄ and W cc̄)

– Use QCD matrix element for heavy flavour content ∗ Use HERWIG and calculation by M. Mangano ∗ CDF assigns a 100% uncertainty to rate

– Use known tagging efficiencies

– Calibrate against W + 1 and W + 2 jet data

Method I predicts larger background than Method II

• Use Method II as more reliable background estimate

Results for SVX Background Calculation

Can compare CDF results for Run 1B

• Summarise major sources of background below

• Estimates for CDF Run 1B data only

• Note that W c background is added to both

W + 1 jet  W + 2 jet  W + 3 jet  W + ≥ 4 jet
Method 1  59 ± 7  20 ± 2  5.0 ± 0.7  1.6 ± 0.2
Fakes  10 ± 2  3.5 ± 0.7  0.9 ± 0.2  0.4 ± 0.1
W bb̄ W cc̄  11 ± 9  5.4 ± 4.3  1.4 ± 1.1  0.4 ± 0.3
W c  12 ± 4  3.3 ± 1.0  0.7 ± 0.3  0.2 ± 0.1
Method 2  38 ± 10  15 ± 5  3.7 ± 1.2  1.1 ± 0.3